what is the value of x in the equation 4(2x+1)=27+3(2x-5)
Answer by checkley79(3341)

4(2x+1)=27+3(2x-5)
8X+4=27+6X-15
8X-6X=27-15-4
2X=8
X=8/2
X=4 ANS.
PROOF:
4(2*4+1)=27+3(2*4-5)
4(8+1)=27+3(8-5)
4*9=27+3*3
36=27+9
36=36